VIBRATIONAL DEPENDENCE OF EXCITED STATE INTRAMOLECULAR PROTON TRANSFER IN 2-(2′-PYRIDYL)PYRROLE IN THE GAS PHASE VIA HIGH RESOLUTION ELECTRONIC SPEC-TROSCOPY. ^{a}^{b}

\underline{PHILIP J. MORGAN}, ADAM J. FLEISHER and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; MICHAL KIJAK and JACEK WALUK, Institute of Physical Chemistry, Polish Academy of Sciences, Kasprzaka 44/52, 01-224 Warsaw, Poland.

Rotationally resolved fluorescence excitation spectra of the S_{1} ←S_{0} origin band and +144 cm^{-1} vibrational band transitions of 2-(2′-Pyridyl)pyrrole (2PP) have been recorded in the collision free environment of a molecular beam. Analyses of these data provide new information about the changes in geometry that occur when 2PP absorbs light. Additionally, significant line broadening is observed in both spectra, which we attribute to an excited state intramolecular proton transfer (ESIPT) reaction. The dynamics and vibrational mode dependence of ESIPT in 2PP will be discussed.

aWork supported by NSF (CHE-0911117)

bSynthesis by Randolph P. Thummel, University of Houston, Houston, Texas 77204-5003